# Axion Dark Matter Archaeology with Primordial Gravitational Waves

### v1

@@ Abstract
We investigate the complementary information to be gained from inflationary gravitational wave (IGW) signals and searches for QCD axion dark matter. We focus on post-inflationary Peccei-Quinn (PQ) breaking axion models that are cosmologically safe. Recent work has shown that a greater number of such models exist. This is because the heavy quarks required for the colour anomaly can provoke a period of heavy quark domination (HQD), which, through decay, dilutes the axion abundance. In this work we show for the first time that the axion dark matter mass can be as low as $m_{a}\sim 10^{-8}\,{\rm eV}$ for models where the heavy quarks decay via dimension 6 terms. This is achieved by allowing the mass of the heavy quarks to differ from the axion decay constant, $m_{Q}\neq f_{a}$. Consequently, the observables that would distinguish between pre- and post-inflationary PQ breaking, $m_{a}$ and the additional relativistic degrees of freedom $\Delta N_{\rm eff}$, now become indiscernible. To solve this, we propose using blue-tilted IGWs to probe HQD. By leveraging the features of the GW signal, future interferometers can probe $m_{Q}$ and $f_{a}$ complementing the sensitivity of haloscope experiments, potentially pinning down all relevant parameter that describe the physics. Specifically, we find that BBO and ET GW detectors will both be able to optimistically probe $f_{a}\gtrsim 10^{14}\,{\rm GeV}$.
1 Introduction
@@
The QCD axion provides a compelling an elegant explanation of both the strong CP problem and dark matter simultaneously. The axion itself is the proposed pseudo-Nambu-Goldstone boson of a new spontaneously broken $U(1)_{\rm PQ}$ symmetry [1, 2, 3, 4]. As a consequence the axion’s mass, $m_{a}$, can naturally be very light and can be related to the breaking scale $f_{a}$ via
$m_{a}=5.7\times 10^{-5}\,{\rm eV}\left(\frac{10^{11}\,{\rm GeV}}{f_{a}}\right).$ (1.1)
@@
Such a light boson in an expanding Universe will undergo a non-thermal production mechanism known as vacuum misalignment [5, 6, 7] (see also reviews in Refs. [8, 9]). This mechanism produces a cold dark matter abundance of axions which can constitute all of the relic dark matter inferred from cosmology and astrophysics [10]. This implies a very high symmetry breaking scale, $f_{a}\sim 10^{11}-10^{12}\,{\rm GeV}$ with the corresponding axion mass, $m_{a}$ being $\mathcal{O}(10)\,\mu{\rm eV}$. Therefore, by searching for evidence of the axion, we will be indirectly probing physics at energy scales far beyond the reach of current particle physics colliders.
@@
Fortunately, the solution to the CP problem necessitates some interactions between the axion and the Standard Model (SM), the strength of which is also suppressed by $f_{a}$. Despite this large suppression, axion emission from supernovae [11] as well as precision measurement of hadron decays at experiments like NA62 [12] and KOTO [13] provide lower bounds on $f_{a}$ [14, 15]. Furthermore, dedicated axion dark matter experiments such as ADMX [16, 17], CAPP [18], UF [19], TASEH [20], HAYSTAC [21] and QUAX [22, 23] have achieved excellent sensitivities, and future experiments like MADMAX [24, 25], ORGAN [26, 27], FLASH [28, 29], BabyIAXO-RADES [30], DMRadio [31] promise to survey large portions of the parameter range where QCD axion dark matter is predicted.
@@
By probing energies far beyond the ${\rm TeV}$-scale we are indirectly exploring physics of the very early Universe, an epoch we know very little about [32]. Indeed we can only say with certainty that there was a period of cosmic inflation which ended and by the time of Big Bang Nucleosynthesis (BBN), $T\sim 3\,{\rm MeV}$, the Universe was dominated by SM radiation. This provides ambiguity about the dynamics of any high-scale extension to the SM such as the those responsible for the PQ mechanism. For example, the consequences of symmetry breaking above or below the temperature of reheating after inflation are different. If the $U(1)_{\rm PQ}$ symmetry is broken after inflation many axion models predict the formation of stable topological defects such as domain walls [33, 34], which require additional mechanisms to destroy [35, 36, 37]. Recently, it has been shown that a number of models that do not have a domain wall problem [38, 39] will generate a period of early heavy quark domination (HQD) [40, 41] akin to early matter domination (EMD). As a consequence the measurable parameters that would distinguish the pre and post-inflationary symmetry breaking scenarios are equivalent. The task is then to find additional probes on the time of $U(1)_{\rm PQ}$ breaking in relation to inflation. We show that inflationary first-order gravitational waves (GWs), in tandem with axion haloscope experiments will be able do this. We promote such tests of axion dark matter in complementary searches between laboratory and GW experiments.
@@
In this work we focus on GWs that are generated from quantum tensor fluctuations of the metric during inflation [42, 43, 44, 45]. The spectrum of such GWs follow a simple power law during radiation domination acting as a tracer of cosmic expansion  [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Any departure from the standard thermal history, such as HQD provides distinctive GW spectral features  [59, 60, 61, 62, 63, 64, 65, 66]. Investigating such features reveals both the time at which early matter domination began and its duration. We will show that this allows us to get a handle on the mass of the particle responsible for HQD, and the rate of their decay. Only when combining this information with axion haloscope experiments can we determine the breaking scale $f_{a}$. We therefore show that complementarity between GW and axion experiments will be vital for reconstructing the relevant particle physics parameters.
@@
This paper is organized as follows, in section 2 we describe effective axion model framework that we utilize in this article. We discuss how EMD is generated and show how such a scenario changes the expected value for $m_{a}$. We then discuss how this complicates phenomenological implications for pre and post-inflationary $U(1)_{\rm PQ}$ scenario. In section 3 we review the modelling of inflationary first-order GWs and how they can be used to detect a period of EMD before BBN. We also describe our limit setting methodology which we undertake for all relevant current and future GW detectors. In section 4 we present our results, combining axion haloscope and GW experiments, we discuss the current status and future potential. Finally we conclude in section 5.
2 Preferred Axion models with Heavy Quarks
@@
The cornerstone of the Peccei-Quinn (PQ) solution [1, 2] to the strong CP (charge parity) problem is to propose a dimension-5 term, the colour anomaly, in addition to the CP violating term allowed by SM gauge invariance. These enter the Lagrangian as
@@
$\mathcal{L}_{\text{PQ}}\supset N\frac{a\left(t,x\right)}{f_{a}}\frac{g_{s}^{2}}{32\pi^{2}}\tilde{G}^{\mu\nu}_{b}G_{\mu\nu}+\theta\frac{g_{s}^{2}}{32\pi^{2}}\tilde{G}^{\mu\nu}_{b}G_{\mu\nu},$ (2.1)
@@
where $a$ is the axion field, $G$ and $\tilde{G}$ are the field strength tensors of the gluon field and its dual respectively. The coefficients $g_{s}$, $\theta$, $N$, and $f_{a}$ are the SM strong coupling constant, the QCD vacuum angle, a model-dependent anomaly coefficient, and the axion decay constant respectively. The first term, under the influence of the QCD potential is dynamically driven to cancel the second term, resulting in no detectable CP violation from strong interactions. How this dimension-5 operator is generated is a model building exercise. Typically there are two approaches: either with SM particles charged under the new $U(1)_{\rm PQ}$ (DSFZ-type [67, 68, 69]) or not (KSVZ-type [70, 71]). In both cases new particles are required, for KSVZ, the minimal requirement is to introduce two new fields: a new $SU(3)_{c}$ charged fermion, $Q$, a heavy quark; and a complex scalar field $\Phi$ which drives the PQ breaking. The PQ Lagrangian is given by
@@
Where $y_{Q}$ is the BSM Yukawa coupling which generates Q mass when PQ symmetry is broken. The PQ charge of $\Phi$ is $1$ and the chiral left-handed $Q_{L}$ and right-handed $Q_{R}$ fields have different charges $\chi_{L}$ and $\chi_{R}$ respectively, but with $|\chi_{L}-\chi_{R}|=1$ such that the above Lagrangian is symmetric under $U(1)_{\rm PQ}$. In the broken phase, the Yukawa term reduces to a mass term with $m_{Q}=y_{Q}f_{a}/\sqrt{2}$. If the breaking scale is below the reheating temperature of the early Universe, $T_{\rm PQ}<T_{\rm RH}$, then the strongly charged $Q$ particles will initially be in thermal equilibrium with the SM plasma. After PQ breaking, the heavy quarks acquire a mass and will freeze-out. As pointed out in Refs. [38, 39], performing standard thermal production would generate a large relic abundance of heavy quarks, violating constraints from cosmology unless they decay quickly enough.
@@
Furthermore, the misalignment mechanism generates a cold axion abundance [5, 6, 7, 72]:
$\Omega_{a}h^{2}\approx 0.12\,\left(\frac{\theta_{\rm i}}{2.15}\right)^{2}\left(\frac{28~{}\mu{\rm eV}}{m_{a}}\right)^{7/6},$ (2.3)
@@
where $\theta_{i}$ is the initial angle of the field, $\theta=a/f_{a}$, which for the $U(1)_{\rm PQ}$ breaking after inflation is approximated to be $\theta_{i}\approx 2.15$. This number is obtained by accounting for multiple Hubble patches undergoing misalignment with an ensemble of $\theta_{i}$ values [73]. This then limits the axion mass to be $m_{a}\,\gtrsim 15\,\mu{\rm eV}$ in order to avoid axion overproduction.
@@
Combining this with the need for fast $Q$ decays, authors of Refs. [38, 39] utilized an effective field theory approach as a guide on what axion models are compatible with standard cosmology, heavy quark decay terms are described by
$\mathcal{L}_{Qq}=\mathcal{L}_{Qq}^{d\leq 4}+\mathcal{L}_{Qq}^{d>4}=\mathcal{L}_{Qq}^{d\leq 4}+\frac{1}{\Lambda^{(d-4)}}\mathcal{O}^{d>4}+{\rm h.c.}\,.$ (2.4)
@@
where $d$ is the mass dimension of the operators. Taking the cut-off scale as the Planck scale $\Lambda=m_{\rm Pl}=1.22\times 10^{19}\,{\rm GeV}$, it was found that the maximum dimension for decay while avoiding the overproduction of axions was $d=5$. However, recently Ref. [40] found that for decays $d\geq 5$, the thermally produced heavy quarks will provoke a period of heavy quark domination (HQD), altering the misalignment mechanism in such a way as to allow for smaller axion masses and larger dimension for $Q$ decay. A period of early matter domination altering the dynamics of the misalignment mechanism has been explored before [74, 75, 76, 77, 78, 79, 80, 81, 82, 83], typically by introducing additional matter content as new physics. Ref. [40] on the other hand pointed out that for KSVZ axion models, the additional matter content is already present. This result means that a greater number of axion models are viable, in Ref. [41] authors enumerate such models. Notably, they found that there are four $d=6$ and two $d=9$ SM charge assignment for the heavy quark that do not suffer from a domain wall problem, this extends the previously known two, KSVZ-I and KSVZ-II, to eight.
2.1 Condition for heavy quark domination

### v2

@@ Abstract
We investigate the complementary information to be gained from inflationary gravitational wave (IGW) signals and searches for QCD axion dark matter. We focus on post-inflationary Peccei-Quinn (PQ) breaking axion models that are cosmologically safe. Recent work has shown that a greater number of such models exist. This is because the heavy quarks required for the colour anomaly can provoke a period of heavy quark domination (HQD), which, through decay, dilutes the axion abundance. In this work we show for the first time that the axion dark matter mass can be as low as $m_{a}\sim 10^{-8}\,{\rm eV}$ for models where the heavy quarks decay via dimension 6 terms. This is achieved by allowing the mass of the heavy quarks to differ from the axion decay constant, $m_{Q}\neq f_{a}$. Consequently, the observables that would distinguish between pre- and post-inflationary PQ breaking, $m_{a}$ and the additional relativistic degrees of freedom $\Delta N_{\rm eff}$, now become indiscernible. To solve this, we propose using blue-tilted IGWs to probe HQD. In scenarios where such a blue tilt is present, the enhanced GW signal allows future interferometers to place non-trivial constraints on the parameters $m_{Q}$ and $f_{a}$, thereby complementing haloscope searches. While some degeneracies with other parameters such as $m_{Q}$ remain, detectors such as BBO and ET will be able to optimistically probe $f_{a}\gtrsim 10^{14}\,{\rm GeV}$.
I Introduction
@@
The QCD axion provides a compelling an elegant explanation of both the strong CP problem and dark matter simultaneously. The axion itself is the proposed pseudo-Nambu-Goldstone boson of a new spontaneously broken $U(1)_{\rm PQ}$ symmetry Peccei and Quinn (1977a, b); Weinberg (1978); Wilczek (1978). As a consequence the axion’s mass, $m_{a}$, can naturally be very light and can be related to the breaking scale $f_{a}$ via
$m_{a}=5.7\times 10^{-5}\,{\rm eV}\left(\frac{10^{11}\,{\rm GeV}}{f_{a}}\right).$ (1)
@@
Such a light boson in an expanding Universe will undergo a non-thermal production mechanism known as vacuum misalignment Preskill et al. (1983); Dine and Fischler (1983); Abbott and Sikivie (1983) (see also reviews in Refs. Sikivie (2008); Marsh (2016)). This mechanism produces a cold dark matter abundance of axions which can constitute all of the relic dark matter inferred from cosmology and astrophysics Cirelli et al. (2024). This implies a very high symmetry breaking scale, $f_{a}\sim 10^{11}-10^{12}\,{\rm GeV}$ with the corresponding axion mass, $m_{a}$ being $\mathcal{O}(10)\,\mu{\rm eV}$. Therefore, by searching for evidence of the axion, we will be indirectly probing physics at energy scales far beyond the reach of current particle physics colliders.
@@
Fortunately, the solution to the CP problem necessitates some interactions between the axion and the Standard Model (SM), the strength of which is also suppressed by $f_{a}$. Despite this large suppression, axion emission from supernovae Carenza et al. (2019) as well as precision measurement of hadron decays at experiments like NA62 Cortina Gil et al. (2017) and KOTO Yamanaka (2012) provide lower bounds on $f_{a}$ Martin Camalich et al. (2020); Alonso-Álvarez et al. (2023). Furthermore, dedicated axion dark matter experiments such as ADMX Asztalos et al. (2010); Braine et al. (2020), CAPP Kim et al. (2024a), UF Hagmann et al. (1996), TASEH Chang et al. (2022), HAYSTAC Zhong et al. (2018) and QUAX Di Vora et al. (2023); Rettaroli et al. (2024) have achieved excellent sensitivities, and future experiments like MADMAX Brun et al. (2019); Egge et al. (2024), ORGAN McAllister et al. (2017); McAllister and Tobar (2020), FLASH Alesini et al. (2017, 2023), BabyIAXO-RADES Ahyoune et al. (2023), DMRadio Brouwer et al. (2022) promise to survey large portions of the parameter range where QCD axion dark matter is predicted.
@@
By probing energies far beyond the $\rm TeV$}-scale we are indirectly exploring physics of the very early Universe, an epoch we know very little about Allahverdi et al. (2020). Indeed we can only say with certainty that there was a period of cosmic inflation which ended and by the time of Big Bang Nucleosynthesis (BBN), $T\sim 3\,{\rm MeV}$, the Universe was dominated by SM radiation. This provides ambiguity about the dynamics of any high-scale extension to the SM such as the those responsible for the PQ mechanism. For example, the consequences of symmetry breaking above or below the temperature of reheating after inflation are different. If the $U(1)_{\rm PQ}$ symmetry is broken after inflation many axion models predict the formation of stable topological defects such as domain walls Zeldovich et al. (1974); Sikivie (1982), which require additional mechanisms to destroy Barr and Kim (2014); Reig (2019); Caputo and Reig (2019). Recently, it has been shown that a number of models that do not have a domain wall problem Di Luzio et al. (2017a, b) will generate a period of early heavy quark domination (HQD) Cheek et al. (2024); Di Luzio et al. (2024) akin to early matter domination (EMD). As a consequence the measurable parameters that would distinguish the pre and post-inflationary symmetry breaking scenarios are equivalent. The task is then to find additional probes on the time of $U(1)_{\rm PQ}$ breaking in relation to inflation. We show that inflationary first-order gravitational waves (GWs), in tandem with axion haloscope experiments will be able do this. We promote such tests of axion dark matter in complementary searches between laboratory and GW experiments.
@@
In this work we focus on GWs that are generated from quantum tensor fluctuations of the metric during inflation Grishchuk (1974); Starobinsky (1979); Rubakov et al. (1982); Guzzetti et al. (2016). The spectrum of such GWs follow a simple power law during radiation domination acting as a tracer of cosmic expansion  Seto and Yokoyama (2003); Boyle and Steinhardt (2008); Boyle and Buonanno (2008); Kuroyanagi et al. (2009); Nakayama and Yokoyama (2010); Kuroyanagi et al. (2013); Jinno et al. (2014); Saikawa and Shirai (2018); Chen et al. (2024); Ghoshal et al. (2024); Haque et al. (2021); Chakraborty et al. (2023); Maity and Haque (2024). Any departure from the standard thermal history, such as HQD provides distinctive GW spectral features  Berbig and Ghoshal (2023); Borboruah et al. (2024a, b); Bernal et al. (2020); Datta and Samanta (2022, 2023); Chianese et al. (2024); Borboruah et al. (2025). Investigating such features reveals both the time at which early matter domination began and its duration. We will show that this allows us to get a handle on the mass of the particle responsible for HQD, and the rate of their decay. Only when combining this information with axion haloscope experiments can we determine the breaking scale $f_{a}$. We therefore show that complementarity between GW and axion experiments will be vital for reconstructing the relevant particle physics parameters.
@@
This paper is organized as follows, in section II we describe effective axion model framework that we utilize in this article. We discuss how EMD is generated and show how such a scenario changes the expected value for $m_{a}$. We then discuss how this complicates phenomenological implications for pre and post-inflationary $U(1)_{\rm PQ}$ scenario. In section III we review the modelling of inflationary first-order GWs and how they can be used to detect a period of EMD before BBN. We also describe our limit setting methodology which we undertake for all relevant current and future GW detectors. In section IV we present our results, combining axion haloscope and GW experiments, we discuss the current status and future potential. Finally we conclude in section V.
II Preferred Axion models with Heavy Quarks
@@
The cornerstone of the Peccei-Quinn (PQ) solution Peccei and Quinn (1977a, b) to the strong CP (charge parity) problem is to propose a dimension-5 term, the colour anomaly, in addition to the CP violating term allowed by SM gauge invariance. These enter the Lagrangian as
@@
$\mathcal{L}_{\text{PQ}}\supset N\frac{a\left(t,x\right)}{f_{a}}\frac{g_{s}^{2}}{32\pi^{2}}\tilde{G}^{\mu\nu}_{b}G_{\mu\nu}+\theta\frac{g_{s}^{2}}{32\pi^{2}}\tilde{G}^{\mu\nu}_{b}G_{\mu\nu},$ (2)
@@
where $a$ is the axion field, $G$ and $\tilde{G}$ are the field strength tensors of the gluon field and its dual respectively. The coefficients $g_{s}$, $\theta$, $N$, and $f_{a}$ are the SM strong coupling constant, the QCD vacuum angle, a model-dependent anomaly coefficient, and the axion decay constant respectively. The first term, under the influence of the QCD potential is dynamically driven to cancel the second term, resulting in no detectable CP violation from strong interactions. How this dimension-5 operator is generated is a model building exercise. Typically there are two approaches: either with SM particles charged under the new $U(1)_{\rm PQ}$ (DSFZ-type Zhitnitsky (1980); Dine et al. (1981); Kim and Carosi (2010)) or not (KSVZ-type Kim (1979); Shifman et al. (1980)). In both cases new particles are required, for KSVZ, the minimal requirement is to introduce two new fields: a new $SU(3)_{c}$ charged fermion, $Q$, a heavy quark; and a complex scalar field $\Phi$ which drives the PQ breaking. The PQ Lagrangian is given by
@@
Where $y_{Q}$ is the BSM Yukawa coupling which generates Q mass when PQ symmetry is broken. The PQ charge of $\Phi$ is $1$ and the chiral left-handed $Q_{L}$ and right-handed $Q_{R}$ fields have different charges $\chi_{L}$ and $\chi_{R}$ respectively, but with $|\chi_{L}-\chi_{R}|=1$ such that the above Lagrangian is symmetric under $U(1)_{\rm PQ}$. In the broken phase, the Yukawa term reduces to a mass term with $m_{Q}=y_{Q}f_{a}/\sqrt{2}$. If the breaking scale is below the reheating temperature of the early Universe, $T_{\rm PQ}<T_{\rm RH}$, then the strongly charged $Q$ particles will initially be in thermal equilibrium with the SM plasma. After PQ breaking, the heavy quarks acquire a mass and will freeze-out. As pointed out in Refs. Di Luzio et al. (2017a, b), performing standard thermal production would generate a large relic abundance of heavy quarks, violating constraints from cosmology unless they decay quickly enough. Whether or not $Q$-decay is fast is determined by their SM and PQ charges. For example, the original KSVZ model has a charge configuration such that the heavy quark is stable Kim (1979). Finding KSZV-type models that enable these heavy quarks to decay fast enough then becomes a model building exercise.
@@
Furthermore, the misalignment mechanism generates a cold axion abundance Preskill et al. (1983); Dine and Fischler (1983); Abbott and Sikivie (1983); Di Luzio et al. (2020):
$\Omega_{a}h^{2}\approx 0.12\,\left(\frac{\theta_{\rm i}}{2.15}\right)^{2}\left(\frac{28~\mu{\rm eV}}{m_{a}}\right)^{7/6},$ (4)
@@
where $\theta_{i}$ is the initial angle of the field, $\theta=a/f_{a}$, which for the $U(1)_{\rm PQ}$ breaking after inflation is approximated to be $\theta_{i}\approx 2.15$. This number is obtained by accounting for multiple Hubble patches undergoing misalignment with an ensemble of $\theta_{i}$ values Grilli di Cortona et al. (2016). This then limits the axion mass to be $m_{a}\,\gtrsim 15\,\mu{\rm eV}$ in order to avoid axion overproduction.
@@
Combining this with the need for fast $Q$ decays, authors of Refs. Di Luzio et al. (2017a, b) utilized an effective field theory approach as a guide on what axion models are compatible with standard cosmology, heavy quark decay terms are described by
$\mathcal{L}_{Qq}=\mathcal{L}_{Qq}^{d\leq 4}+\mathcal{L}_{Qq}^{d>4}=\mathcal{L}_{Qq}^{d\leq 4}+\frac{1}{\Lambda^{(d-4)}}\mathcal{O}^{d>4}+{\rm H.c.}\,.$ (5)
@@
where $d$ is the mass dimension of the operators. Taking the cut-off scale as the Planck scale $\Lambda=m_{\rm Pl}=1.22\times 10^{19}\,{\rm GeV}$, it was found that the maximum dimension for decay while avoiding the overproduction of axions was $d=5$. This would then exclude models with $Q$ charges with $d\geq 6$ decays at lowest order, such as the model where $Q$ has the SM charges $\left(SU(3)_{\rm c},SU(2)_{\rm W},U(1)_{Y}\right)=(3,1,-7/3)$, which at lowest order has the decay operator $\mathcal{O}=\bar{Q}_{L}d_{R}\left(\bar{e}_{R}^{c}\,e_{R}\right)$. However, recently Ref. Cheek et al. (2024) found that for decays $d\geq 5$, the thermally produced heavy quarks will provoke a period of heavy quark domination (HQD), altering the misalignment mechanism in such a way as to allow for smaller axion masses and larger dimension for $Q$ decay. A period of early matter domination altering the dynamics of the misalignment mechanism has been explored before Steinhardt and Turner (1983); Lazarides et al. (1990); Kawasaki et al. (1996); Visinelli and Gondolo (2010); Nelson and Xiao (2018); Ramberg and Visinelli (2019); Arias et al. (2021); Xu (2023); Arias et al. (2023, 2024); Mazde and Visinelli (2023), typically by introducing additional matter content as new physics. Ref. Cheek et al. (2024) on the other hand pointed out that for KSVZ axion models, the additional matter content is already present. This result means that a greater number of axion models are viable, in Ref. Di Luzio et al. (2024) authors enumerate such models. Notably, they found that there are four $d=6$ and two $d=9$ SM charge assignment for the heavy quark that do not suffer from a domain wall problem, this extends the previously known two, KSVZ-I and KSVZ-II, to eight.
II.1 Condition for heavy quark domination